Tourists covered 640 km for a 4-hour ride and a 7-hour ride by train. What is the speed of the train, if it is 5 km greater than the speed of the car?
Answer by Boreal(15235)

ride speed is x; distance is 4x
train speed is x+5; distance is 7x+35
The two add to 640
11x+35=640
11x=605
x=55 km/h ride
x+5=60 km/h train
55*4=220
60*7=420
They add up to 640 km.